OpenStudy (anonymous):

How would i graph this? f(x) = 3(1.5)x

OpenStudy (anonymous):

@amistre64 @Algorithmic

OpenStudy (michele_laino):

please note that: \[1.5=\frac{ 3 }{ 2 }\] so your function can be rewritten as below: \[f(x)=\frac{ 9 }{ 2 }x\] which is a line passing through the origin and with a slope of ? please continue

OpenStudy (anonymous):

@Michele_Laino Im so completely lost. I know that we would use Rise/Run.

OpenStudy (michele_laino):

please, your function is f(x)=(9/2)x, do you agree?

OpenStudy (anonymous):

Yes, With you so far.

OpenStudy (triciaal):